A 45 foot speaker wire is to be cut into three pieces. One piece is to be 15 feet long. Of the remaining pieces, one must be 2 feet less than 3 times the length of the other. Find the length of the shorter piece.
Answer by mananth(16946)

total length = 45
..
first piece 15 feet
let second be x
third piece =30-x
...
30-x=3x-2
+x
30=3x+x-2
30=4x-2
+2
30+2=4x
32=4x
/4
8=x
The lengths are 8, 22, & 15